OpenStudy (anonymous):

Calculus, multiple choice: Find the absolute extrema of f(x)= (x^3 -4x^2 +7x)/x on the interval (0,3] Help is very much appreciated!

OpenStudy (anonymous):

maximum: none, minimum: 4 maximum: 7, minimum: 4 maximum: 7, minimum: 3 maximum: none, minimum: 3 none of these

OpenStudy (anonymous):

these are my choices. I think the fact that this is not a completely closed interval is getting to me. how should I start?

OpenStudy (anonymous):

i'll just use the quotient rule to find derivative and then see where that goes... give me a sec.

OpenStudy (anonymous):

derivative simplifies to f'= 2x-4 ?

OpenStudy (phi):

yes. It would have been ok to simplify the original equation to x^2 -4x + 7 and f' = 2x -4 = 0

OpenStudy (anonymous):

ok I went around my elbow. :3 so f'=0 when x=2... ah, what do I do next?

OpenStudy (anonymous):

plug it into f ? f(2)= 3

OpenStudy (phi):

one way to go is evaluate the original function at x=2 compare to the value of the function at x=3 (and at x=0, though 0 is not in the domain... it gives us an idea of the shape of the curve)

OpenStudy (anonymous):

concave upward parabola in upper two quadrants is what I am getting

OpenStudy (anonymous):

so there would be no maximum, only a minimum

OpenStudy (phi):

sounds correct. the value at x=3 is 4 the value at x=0 is 7 (but we never reach 7 because 0 is excluded from the domain.) so only a min exists in the interval (0,3]

OpenStudy (anonymous):

so, with these answer choices would it be A? because the minimum is at y=4?

OpenStudy (anonymous):

we can cross out B and C easily, and E too, I believe... A vs. D is tricky and I'm not entirely sure which one it is asking for

OpenStudy (phi):

what happened to plug it into f ? f(2)= 3

OpenStudy (anonymous):

what? oh, sorry I was confused by your mention of (3,4) being a point. so (2,3) is a point, too. so the minimum is y= 3 and the answer is D

OpenStudy (phi):

yes, no max, min of 3 (at x=2)

OpenStudy (anonymous):

ok thank you very much! I had just been worried that the interval was a 'hybrid' of sorts, but now I know that you can solve the same way! :) you get a medal, and i'll close the question. again, thanks for your time!